# GATE OpenING: A Comprehensive Benchmark for Judging Open-ended Interleaved Image-Text Generation

### v1

@@ Appendix D Extension of OpenING
The results highlight that MiniGPT-5OpenING achieves significant improvements over the baseline MiniGPT-5 models across all metrics. For example, in the Without Tie (w/o Tie) scenario, the finetuned model shows a substantial 37.39% relative improvement over the MiniGPT-5 baseline. These findings confirm that training on a specialized interleaved image-text dataset such as the Dev Set of OpenING enhance the model with better contextual understanding and alignment capabilities for generating coherent interleaved image-text content. Further studies are ongoing to improve the performance of SOTA models.
Appendix E Limitations of This Study
@@
Although the OpenING benchmark represents a significant step forward in evaluating interleaved image-text generation, it faces several limitations that present opportunities for improvement. First, while OpenING expands task diversity with 56 tasks across 23 meta-topics, some real-world scenarios remain underrepresented or oversimplified, potentially limiting the generalizability to practical applications. Tasks requiring fine-grained understanding or multi-step reasoning need to be supplied to capture real-world needs. Second, though the IntJudge model improves alignment with human evaluations, its generalizability is still constrained by the diversity and quality of training data. The benchmark’s reliance on human-annotated data to establish ground truth and train judge models is both labor-intensive and costly. While the proposed Reference-Augmented Generation (RAG) approach helps scale training data, manual annotations remain a critical component for ensuring quality and alignment with human preferences. Furthermore, the computational resources required for training and deploying IntJudge present scalability challenges, potentially limiting accessibility for researchers with fewer resources.
@@
In addition, current interleaved image-text generation methods still struggle with producing high-quality, coherent interleaved content, particularly in multi-step tasks that require maintaining consistency across generated images and text. Issues like content incoherence, poor image quality, and mismatches between generated text and images persist across evaluated models, particularly in end-to-end approaches. To tackle these issues, more advanced MLLMs trained with a large-scale interleaved image-text dataset are to be investigated. What is more, building a sufficiently comprehensive, diverse, and representative dataset is expected to greatly promote the development of multimodal generation. These limitations underscore the need for continued development of more diverse datasets and more robust evaluation frameworks to address the complexities of interleaved generation evaluation, enabling more practical interleaved image-text generation methods and pushing forward the boundary of future MLLMs.
Figure 18: Representative examples from 23 Meta-Topics in Our OpenING Benchmark.
Figure 24: Illustration of 23 pairs of generated outputs across 23 meta-topics. The gold medal represents the winner of pairwise comparison and the silver medal denotes a more favorable output in a tie scene of the pairwise comparison. The meta-topic name and selected task name for each pairwise data instance are given, which are separated by colons.

### v3

@@ Appendix E Finetuning MLLMs on OpenING
The results highlight that MiniGPT-5OpenING achieves significant improvements over the baseline MiniGPT-5 models across all metrics. For example, in the Without Tie (w/o Tie) scenario, the finetuned model shows a substantial 37.39% relative improvement over the MiniGPT-5 baseline. These findings confirm that training on a specialized interleaved image-text dataset such as the Dev Set of OpenING enhances the model with better contextual understanding and alignment capabilities for generating coherent interleaved image-text content. Further studies are ongoing to improve the performance of SOTA models.
Appendix F Limitations of This Study
@@
Although OpenING advances interleaved image-text generation evaluation, several limitations remain for improvement. First, despite covering diverse tasks (56 tasks across 23 meta-topics), some real-world scenarios remain underrepresented or oversimplified, potentially limiting the generalizability to practical applications. Tasks requiring fine-grained understanding or multi-step reasoning need to be supplied to capture real-world needs. Second, although the IntJudge model improves alignment with human evaluations, its generalizability remains limited by the diversity and quality of training data. The benchmark’s reliance on human-annotated data to establish ground truth and train judge models is labor-intensive and expensive. While our Reference-Augmented Generation (RAG) approach helps scale training data, manual annotations remain a critical component for ensuring quality and alignment with human preferences. Furthermore, the computational resources required for training and deploying IntJudge present scalability challenges, potentially limiting accessibility for researchers with fewer resources.
@@
In addition, current interleaved image-text generation methods still struggle with producing high-quality, coherent interleaved content, particularly in multi-step tasks that require maintaining consistency across generated images and text. Issues like content incoherence, poor image quality, and mismatches between generated text and images persist across evaluated models, particularly in end-to-end approaches. To tackle these issues, more advanced MLLMs trained with a large-scale interleaved image-text dataset are to be investigated. Moreover, while our IntJudge demonstrates significant advantages over GPT-based evaluators, we recognize two aspects for improvement: 1) Potential biases may rise from crowdsourced data (e.g., human preferences in aesthetic judgment), and 2) The current foundation model of IntJudge mainly supports English and Chinese. Building a sufficiently comprehensive, diverse, and representative dataset is expected to greatly promote the development of multimodal generation.
These limitations underscore the need for continued development of more diverse datasets and more robust evaluation frameworks to address the complexities of interleaved generation evaluation, enabling more practical interleaved image-text generation methods and pushing forward the boundary of future MLLMs. Future work may benefit from diversifying data sources and cross-cultural annotations, expanding multilingual capabilities, and implementing debiasing techniques to improve fairness.
Acknowledgement
This work is partially supported by the National Key R&D Program of China (No. 2022ZD0160101 and No. 2022ZD0160102).